PHOTODISSOCIATION DYNAMICS OF A TRIATOMIC PSEUDO-DIHALIDE: ABSORPTION CROSS SECTION AND DYNAMICS OF SOLVATED ICN^{−}

\underline{JOSHUA P. MARTIN}, QUANLI GU^{a}, JOSHUA P. DARR^{b}, JILA, Department of Chemistry and Biochemistry University of Colorado at Boulder, Boulder, CO 80309; ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210; and W. CARL LINEBERGER, JILA, Department of Chemistry and Biochemistry, University of Colorado at Boulder, Boulder, CO 80309.

We report the photoabsorption cross section and photoproduct branching ratios of mass-selected bare ICN^{−} and ICN^{−}(CO_{2}) following excitation to the A′ ^{2}Π_{1/2} electronic excited state. Previous studies of CO_{2} solvated-heteronuclear dihalides, IX^{−}(CO_{2})_{n} (X=Cl, Br), reported three excited state selective classes of photoproducts: I^{−}, X^{−}, and IX^{−} based clusters. Photoabsorption of bare ICl^{−} and IBr^{−} that leads to population in the A′ ^{2}Π_{1/2} state have maxima near 680 nm and 740 nm, respectively, and result in I^{−} photoproducts exclusively over the entire band corresponding to A′ ^{2}Π_{1/2} ← X^{2}Σ_{1/2} excitation. Interestingly, following excitation of bare ICN^{−} to the comparable state (430-650 nm, maximum at 490 nm), I^{−} is the dominant ionic photoproduct, but CN^{−} photoproducts are observed as well. When a single CO_{2} solvent molecule is added to ICN^{−}, the same A′ ^{2}Π_{1/2} ← X^{2}Σ_{1/2} excitation results in apparent charge transfer within the complex. Therefore, the observed ionic photoproducts are not just the expected I^{−} and I^{−}(CO_{2}), but CN^{−} and solvated CN^{−}(CO_{2}) photoproducts are also significant products. Analysis of the experimental results using calculated potential energy curves of ICN^{−} reveals intriguing dynamics of the photoexcited triatomic pseudo-dihalide. Supported by NSF and AFOSR.

aPresent address: Department of Chemistry, University of Virginia, Charlottesville, VA 22904

bPresent address: Department of Chemistry, University of Nebraska, Omaha, NE 68182